OpenStudy (anonymous):

What is the solution to the rational equation 5 over x¨Csquared minus 3x plus 2 - 1 over x minus 2 = 1 over 3x minus 3 ?

OpenStudy (anonymous):

let me see if i can write this, and you tell me if it is correct

OpenStudy (anonymous):

\[\frac{5}{x^2-3x+2}-\frac{1}{x}=\frac{1}{3x-3}\]

OpenStudy (anonymous):

maybe this \[\frac{5}{x^2-3x+2}-\frac{1}{x-2}=\frac{1}{3x-3}\]

OpenStudy (anonymous):

maybe?

OpenStudy (anonymous):

yes thats the equation

OpenStudy (anonymous):

ok then we can do it first you have to add on the left hand side

OpenStudy (anonymous):

\[\frac{5}{x^2-3x+2}-\frac{1}{x-2}=\frac{1}{3x-3}\]factor the first denominator \[\frac{5}{(x-1)(x-2)}-\frac{1}{x-2}=\frac{1}{3x-3}\] then least common denominator on the left is \((x-1)(x-2)\) so subtract via \[\frac{5}{(x-1)(x-2)}-\frac{1}{x-2}=\frac{5-(x-1)}{(x-1)(x-2)}\] \[=\frac{6-x}{(x-1)(x-2)}\]

OpenStudy (anonymous):

giving us \[\frac{6-x}{(x-1)(x-2)}=\frac{1}{3x-3}\] then we can "cross multiply" to get \[(3x-3)(6-x)=(x-1)(x-2)\]

OpenStudy (anonymous):

we can even simplify a little before solving because the left hand side now factors as \[3(x-1)(6-x)=(x-1)(x-2)\] cancel the \((x-1)\) which we are allowed to do because that factor cannot be zero, and we get \[3(6-x)=x-2\] \[18-x=x-2\] \[20=2x\] \[10=x\]

OpenStudy (anonymous):

hold on i made a mistake somewhere!

OpenStudy (anonymous):

wow a really bad bush league mistake start with \[3(6-x)=x-2\] \[18-3x=x-2\] \[20=4x\] \[x=5\] now i am embarrassed

OpenStudy (anonymous):

thanks !